OpenStudy (anonymous):

$18,887 is invested, part at 14% and the rest at 8%. If the interest earned from the amount invested at 14% exceeds the interest earned from the amount invested at 8% by $1360.92, how much is invested at each rate? (round to two decimal places if necessary.)

OpenStudy (anonymous):

So what is the amount for 14%??

OpenStudy (amistre64):

the part that is not invested at 8% ..

OpenStudy (anonymous):

sorry misread the question too :(

OpenStudy (anonymous):

x= amount at 14% 18887-x = amount at 8% .14x=.08(18887-x)+1360.92

OpenStudy (amistre64):

a + b = 18,887 a(.14) = b(.08) + 1360

OpenStudy (anonymous):

14x=8(18887-x)+136092 22x=287188 x=13054 other amount 5833

OpenStudy (amistre64):

a + b = t a = t - b a(x) = b(y) + r (t-b)(x) = b(y) + r tx -bx = by +r tx-r = by+bx tx-r = b(y+x) (tx-r)/(y+x) = b 18887(.14) - 1360.92 -------------------- = b .14 + .08

OpenStudy (amistre64):

5833 = b invested at 8% right?

OpenStudy (anonymous):

I am not for sure... So what amount would be invested at 14%?

OpenStudy (anonymous):

13054

OpenStudy (amistre64):

the amount that is not invested at 8% ...

OpenStudy (amistre64):

if you cant subtract by now, there isnt much we can do for you ...